Do (1,1), (2,2), and (3,4) lie on a straight line? And explain
Answer by Cromlix(4381)

Gradient = y2 - y1/x2 - x1
Gradient of the line connecting (1,1) and (2,2)
= 2 - 1/2 - 1 = 1
Gradient of the line connecting (2,2) and (3,4)
= 4 - 2/3 - 2 = 2
Different gradients, therefore not on the same straight line.
Hope this helps.
:-)